OpenStudy (anonymous):

The general equation of a plane is Ax + By + Cz = D, where A, B, C, and D are real numbers and A is nonnegative. Find the equation of the plane containing the points (6, 0, 0), (0, 8, 0), and (0, 0, 3). Show each step of your process. Then graph the plane.

OpenStudy (perl):

solve me the simultaneous equations plug in x = 6, y = 0 , z = 0 , solve for D. similarly for the other points . you will get : 6A = D 8B= D 3C= D A = D/6 B = D/8 C = D / 3

OpenStudy (anonymous):

I suck in math.. Don't know how to do that..

OpenStudy (perl):

D/6*x + D/8*y + D / 3 *z = D factor out D, cancel, then multiply both sides by 6*8*3 (or not)

OpenStudy (anonymous):

Or not? what do you mean by that?

OpenStudy (perl):

you can leave it as 1/6*x + 1/8*y + 1 / 3 *z = 1

OpenStudy (anonymous):

But which would be "more" correct?

OpenStudy (perl):

that depends on style. i like that form, but if you want integer coefficients...

OpenStudy (anonymous):

1/6*x + 1/8*y + 1 / 3 *z = 1 is that the equation?

OpenStudy (perl):

yes

OpenStudy (anonymous):

and what were your steps to find that?

OpenStudy (perl):

the first step, i plugged in x = 6, y = 0 , z = 0 into the given equation, and solved for D.

OpenStudy (perl):

since we were given that (6,0,0) was on the plane

OpenStudy (anonymous):

Do you think you could write out all that you did? I am really stressed about this equation and its due tomorrow morning.. Been trying to figure it out for two weeks now..

OpenStudy (perl):

Given (6,0,0) is on the plane. This means x = 6, y = 0 , z = 0 plug into equation. A(6) + B(0) + C(0) = D A*6 = D A = D/6 thats the first step, now do the same for the next point (0,8,0)

OpenStudy (anonymous):

Oh ok! Could I send you the graph so you can check if its correct?

OpenStudy (perl):

ok

OpenStudy (anonymous):

I dont know how to put a photo in... I dont have it in a folder.. I have it in google drive

OpenStudy (perl):

can you take a screen shot of it

OpenStudy (anonymous):

is it correct?

OpenStudy (perl):



OpenStudy (anonymous):

?

OpenStudy (perl):